# GeoLLaVA-8K: Scaling Remote-Sensing Multimodal Large Language Models to 8K Resolution

### v1

@@ Abstract
Ultra-high-resolution (UHR) remote sensing (RS) imagery offers valuable data for Earth observation but pose challenges for existing multimodal foundation models due to two key bottlenecks: (1) limited availability of UHR training data, and (2) token explosion caused by the large image size. To address data scarcity, we introduce SuperRS-VQA (avg. 8,376$\times$8,376) and HighRS-VQA (avg. 2,000$\times$1,912), the highest-resolution vision-language datasets in RS to date, covering 22 real-world dialogue tasks. To mitigate token explosion, our pilot studies reveal significant redundancy in RS images: crucial information is concentrated in a small subset of object-centric tokens, while pruning background tokens (e.g., ocean or forest) can even improve performance. Motivated by these findings, we propose two strategies: Background Token Pruning and Anchored Token Selection, to reduce the memory footprint while preserving key semantics. Integrating these techniques, we introduce GeoLLaVA-8K, the first RS-focused multimodal large language model capable of handling inputs up to 8K×$\times$8K resolution, built on the LLaVA framework. Trained on SuperRS-VQA and HighRS-VQA, GeoLLaVA-8K sets a new state-of-the-art on the XLRS-Bench. Datasets and code will be released at GeoLLaVA-8K.
@@ 1 Introduction
Can we develop MLLMs tailored for RS data with resolutions far beyond current limits (e.g., up to 8K×$\times$8K)?
@@
In response to this question, we identify two critical challenges that remain underexplored: (1) The lack of image-text training data for UHR RS images; (2) The massive number of visual tokens in UHR RS imagery increases training difficulty for MLLMs. To address the first issue, we introduce two novel multimodal RS datasets featuring large image sizes: SuperRS-VQA (about 8K×$\times$8K) and HighRS-VQA (about 2K×$\times$2K). To our knowledge, they are the largest image-size RS vision-language datasets to date, covering 22 real-world subtasks and significantly surpassing previous RS image-text datasets in both scale and diversity.
For the second problem, from our intuition, the excessive sequence length of visual tokens may lead to two major issues in MLLM’s training: (1) Expensive Computation Overhead: Current models are not designed to operate at such scales. For example, directly adapting LLaVA to 8K×$\times$8K inputs results in memory overflow. (2) Low Semantic Density: UHR RS imagery contains an abundance of homogeneous background tokens, which contribute little useful information. In contrast, semantically rich foreground tokens are sparse and risk being overlooked without effective token management.
@@ 2 Related Work
Remote Sensing Multimodal Datasets. With the rapid progress of large multimodal models in general domains, the RS field has witnessed significant advancements in MLLM development  geochat ; lhrsbot ; rsgpt ; earthgpt ; skysensegpt , leading to the emergence of several RS-specific vision-language datasets. RSVQA rsvqahr comprises image/question/answer triplets, with questions and answers derived from OpenStreetMap (OSM). RSIVQA rsivqa generates samples automatically using existing scene classification and object detection datasets. RSSA h2rsvlm targets hallucination, while FIT-RSRC skysensegpt focuses on object relationship understanding. VRSBench vrsbench includes 29,614 images, 52,472 object references, and 123,221 QA pairs. LRS-VQA lrsvqa constructs a large remote sensing VQA benchmark with 7,333 QA pairs across 8 categories, with image length up to 27,328 pixels. Recent datasets for the SFT stage of MLLM training include GeoChat_Instruct geochat , ChatEarthNet ChatEarthNet , VHM_sft vhm , FIT-RS skysensegpt , and MMRS-1M earthgpt . However, most of these datasets are aggregated from existing sources and contain limited original annotations. Critically, their average image resolution remains below 1K×1K, which is insufficient for real-world UHR RS tasks. Recently, XLRS-Bench (xlrs-bench, ) introduces the largest image-size RS benchmark to date (average 8,500×8,500) for evaluation purposes only, underscoring the persistent lack of corresponding UHR training data in the RS MLLM field.
@@
Multimodal Large Language Model. Leveraging advanced large language models (LLMs) like GPTs (gpt4, ) and LLaMA (llama, ), MLLMs have demonstrated strong visual understanding and reasoning capabilities learning ; DecodingTrust . Proprietary models such as Gemini (gemini, ) and GPT-4o (gpt4, ), along with open-source alternatives like Qwen-VL (qwen, ), InternLM-XComposer (InternLM-XComposer, ), MiniCPM (minicpm, ), LLaVA (visualllama, ), and MiniGPT-4 (minigpt, ), show competitive performance but are typically limited to input resolutions of 2K-4K. To overcome this limitation, LLaVA-Next (llava-next, ) processes images in patches and connects them via global tokens, Monkey (monkey, ) and LLaVA-UHD (llava-uhd, ) compress patches to reduce redundancy, Cambrian (cambrian, ) uses learnable queries for multi-scale interaction, and SliME (slime, ) applies dual compression to preserve global context. In the RS domain, several MLLMs have been developed. For example, GeoChat (geochat, ) enables multi-task RS dialogue based on the LLaVA-1.5 framework (llava15, ). EarthGPT (earthgpt, ) unifies multisensor interpretation tasks by converting RS annotations into question-answer pairs. Meanwhile, LHRS-Bot (lhrsbot, ) improves vision-language understanding through multi-level alignment and curriculum learning.
Nevertheless, both general-domain and RS MLLMs struggle to effectively handle UHR RS imagery. Universal MLLMs lack domain adaptation and are unable to meet the higher resolution requirements of real-world RS applications. Although RS-specific MLLMs incorporate domain knowledge, they still operate at significantly lower resolutions, i.e., less than 1K, highlighting the urgent need for a UHR (8K×8K) and domain-aligned MLLMs in the RS field.
@@ 3 Dataset
To address this, following XLRS-Bench (xlrs-bench, ), we first manually annotated 12K UHR samples. Notably, existing MLLMs such as GPT-4o (gpt4o, ) either encounter memory overflow or generate low-quality outputs when processing UHR RS imagery. Therefore, we adopt a fully manual annotation approach in this work.
Nevertheless, due to the high time and labor costs, scaling manual annotation of UHR data is impractical. To address this, we develop a semi-automated annotation pipeline that leverages GPT-4o (gpt4o, ) alongside existing RS detection and segmentation datasets, generating 100K medium-to-high-resolution (MHR, 2K×2K) image-text pairs. To bridge the distribution gap between MHR and UHR data, we further apply an influence-based data selection method built on the LESS framework xia2024less . Next, we detailed the procedure of constructing datasets.
@@
LLaVA-Next LLaVA-Next-2K_SFT llava-next 738K 512$\times$512 44.6
LLaVA-Next LLaVA-Nex-2K_SFT llava-next 738K 512$\times$512 40.4$\downarrow$
GeoChat_Instruct geochat 283k 632$\times$619
ChatEarthNet ChatEarthNet 6k 256$\times$256
VHM_sft vhm 150k 643$\times$638
FIT-RS skysensegpt 1200k 512$\times$512
MMRS-1M earthgpt 308K 499$\times$491
VRSBench-train vrsbench 85k 512$\times$512
@@
- Multi-image questions 150
- Average question length 128
- Average answer length 22
- Object category 18 (Coarse)
- Object size (pixel ratio) 0.14%
- Average Resolution 8,376$\times$8,378
@@
- Average question length 172
- Average answer length 41
- Object category 60
- Object size (pixel ratio) 1.02%
- Average Resolution 2,000$\times$1,912
@@ 4.1 Overwhelming Background Tokens Hinder MLLM Fine-Tuning on RS Data
In order to learn about a holistic recognition for the background ratio in UHR RS scenes, we conduct preliminary experiments, where the LLaVA-1.5 llava15 is employed. Specifically, we randomly selected 100 images from the XLRS-Bench (xlrs-bench, ), covering natural backgrounds such as sea surfaces, forests, and fields, to provide a diverse set of test samples for the evaluation of background redundancy. Then, we extract 64×64 small images through non-overlapped sliding windows, and sequentially input each into the model.
We evaluated the background rate of the UHR RS scene using two methods, following the previous work (ablation, ): Generative Description: We prompt the model with “Describe the image” and analyze the generated description to extract key information. Binary Polling: We ask the model “Is this image mainly background?” where the background is defined as low-information natural areas (e.g., sea, deserts, vegetation), excluding artificial structures (e.g. buildings, roads, urban areas).
Figure 4: Examples of tokens and their positions that the $logit$ $lens$ yields in the late layers.
@@
To better understand the semantic information of visual tokens, we use the $logit$ $lens$ technique (LogitLens, ). For each layer, we decode the activation at each token position via unembedding. Details are shown in the appendix.
With the $logit$ $lens$ technique, we found that in large background areas, only a small portion of the information in the visual tokens is effectively mapped to specific semantic scenes, while most of the background tokens do not align with clear semantic representations (Fig. 4(b)).
@@
Then, we feed the whole UHR RS imagery into MLLM, where the LLaVA-Next-2K longva , a variant of LLaVA-Next llava-next , is adopted to support higher resolution as far as possible. Initially, we directly perform the inference on XLRS-Bench xlrs-bench , whose images have an average resolution of 8K×8K. Naturally, the resulting long visual token sequences lead to out-of-memory (OOM) errors, aligning with the first challenge presented in Section 1.
@@
To address the OOM issue, we increase the pooling layer size from the default 2×2 to 6×6, allowing LLaVA-Next-2K to process 9K×9K resolution, called LLaVA-Next-9K. To further reduce redundant background information, we use a background-aware token selection strategy. By calculating semantic similarity between visual tokens and typical background terms (e.g., ocean, desert), we score and rank tokens, then discard the top 50% most redundant ones. The process involves: (1) extracting embedding vectors for typical background terms; (2) calculating each token’s highest similarity score with these embeddings as its background score; and (3) ranking tokens by score and removing the top 50%. This approach reduce the background token count. The result is referred to as LLaVA-Next-9K-Pruning.
@@ 4.2 Scarce Object Tokens Drive MLLM Fine-Tuning on RS Data
We conduct ablation experiments to test whether the foreground information is concentrated in specific visual tokens. By removing selected tokens and observing the drop in recognition performance, we assess their importance. The images are sourced from XLRS-Bench, with two pre-processing steps applied to ensure reliability: High signal-to-noise sub-images and Hallucination control. Details are shown in the appendix.
Following this procedure, the final dataset includes 1,189 VQA pairs. Fig. 6 provides an overview of our ablation. Details are shown in the appendix.
@@
Tab. 3 shows that ablating object tokens significantly hinders the model’s ability to recognize targets. Generative decrease and VQA decrease experiments are following the previous work ablation . A larger percentage drop in performance indicates a greater impact of the ablation, meaning the model is more likely to answer incorrectly, thus suggesting that the ablated tokens contain more localized object information. Notably, with a comparable number of ablated tokens, removing object tokens consistently results in a larger performance degradation than random ablation, highlighting the precise localization of object-specific information. $logit$ $lens$ analysis LogitLens further reveals that visual tokens for scarce objects efficiently converge to accurate semantic representations (Fig. 4 (a)). Collectively, these findings underscore a key insight in RS: object tokens not only encode critical information but also align closely with actual visual features.
@@ 4.3 Background Token Pruning and Anchored Token Selection
where $n$ is the number of remaining image tokens, $d$ is the dimension of hidden states, and ${\bm{W}}_{Q},{\bm{W}}_{V}$ are the query and key projection matrices of this encoder layer. Note that we utilize the attention map at the second-to-last layer of the visual encoder ($i.e.$, the output layer of CLIP-ViT in LLaVA-Next (llava-next, )).
Given a compression ratio $r$, we calculate the final number of tokens to retain as $R=n\times r$. Ultimately, we select the top $R$ tokens with the highest attention scores from the attention map ${\bm{a}}_{[\texttt{CLS}]}$, obtaining the final retained tokens $Z_{v}^{\prime}\in\mathbb{R}^{R\times d}$.
@@ 5 Experiments
Experimental Setup. We use XLRS-Bench xlrs-bench for evaluation. The MLLMs evaluated on XLRS-Bench are grouped into three categories: (a) open-source MLLMs; (b) closed-source MLLMs and (c) the specialized RS model. For fair comparison, we used a zero-shot setting with uniform prompts for all MLLMs, including our work. The appendix details the architecture and parameter sizes of each open-source MLLMs, and includes additional results across various settings. Except for GeoChat which was evaluated using its native framework, all other models were evaluated using LMMs-Eval zhang2024lmmsevalrealitycheckevaluation ; lmms_eval2024 .
Following XLRS-Bench xlrs-bench , we evaluated the accuracy and reported of L-1 dimension for the VQA task, with L-3 and L-4 results available in the appendix.
Main Results.
After fine-tuning on SuperRS-VQA and HighRS-VQA, our GeoLLaVA-8K delivers outstanding performance across various evaluation tasks. It not only outperforms domain-specific models but also surpasses all existing open- and closed-source models, including the latest Qwen2.5 and InternVL3. Remarkably, with just 7B parameters, GeoLLaVA-8K even outperforms Qwen2.5-VL-72B, the largest and best-performing open-source MLLMs. This impressive gain stems from the high-quality dataset and the targeted compression strategy designed for the low semantic density of RS imagery.

### v2

@@ Abstract
[…]
@@ 1 Introduction
Can we develop MLLMs tailored for RS data with resolutions far beyond current limits (e.g., up to $\times8K$)?
@@
In response to this question, we identify two critical challenges that remain underexplored: (1) The lack of image-text training data for UHR RS images; (2) The massive number of visual tokens in UHR RS imagery increases training difficulty for MLLMs. To address the first issue, we introduce two novel multimodal RS datasets featuring large image sizes: SuperRS-VQA (about $\times8K$) and HighRS-VQA (about $\times2K$). To our knowledge, they are the largest image-size RS vision-language datasets to date, covering 22 real-world subtasks and significantly surpassing previous RS image-text datasets in both scale and diversity.
For the second problem, from our intuition, the excessive sequence length of visual tokens may lead to two major issues in MLLM’s training: (1) Expensive Computation Overhead: Current models are not designed to operate at such scales. For example, directly adapting LLaVA to $\times8K$ inputs results in memory overflow. (2) Low Semantic Density: UHR RS imagery contains an abundance of homogeneous background tokens, which contribute little useful information. In contrast, semantically rich foreground tokens are sparse and risk being overlooked without effective token management.
@@ 2 Related Work
Remote Sensing Multimodal Datasets. With the rapid progress of large multimodal models in general domains, the RS field has witnessed significant advancements in MLLM development  geochat ; lhrsbot ; rsgpt ; earthgpt ; skysensegpt , leading to the emergence of several RS-specific vision-language datasets. RSVQA rsvqahr comprises image/question/answer triplets, with questions and answers derived from OpenStreetMap (OSM). RSIVQA rsivqa generates samples automatically using existing scene classification and object detection datasets. RSSA h2rsvlm targets hallucination, while FIT-RSRC skysensegpt focuses on object relationship understanding. VRSBench vrsbench includes 29,614 images, 52,472 object references, and 123,221 QA pairs. Recent datasets for the SFT stage of MLLM training include GeoChat_Instruct geochat , ChatEarthNet ChatEarthNet , VHM_sft vhm , FIT-RS skysensegpt , and MMRS-1M earthgpt . However, most of these datasets are aggregated from existing sources and contain limited original annotations. Critically, their average image resolution remains below $\times1K$, which is insufficient for real-world UHR RS tasks.Recently, XLRS-Bench (xlrs-bench, ) introduces the largest image-size RS benchmark to date (average $\times8,500$) for evaluation purposes only, underscoring the persistent lack of corresponding UHR training data in the RS MLLM field.
@@
Multimodal Large Language Model. Leveraging advanced large language models (LLMs) like GPTs (gpt4, ) and LLaMA (llama, ), MLLMs have demonstrated strong visual understanding and reasoning capabilities learning ; DecodingTrust . Proprietary models such as Gemini (gemini, ) and GPT-4o (gpt4, ), along with open-source alternatives like Qwen-VL (qwen, ), InternLM-XComposer (InternLM-XComposer, ), MiniCPM (minicpm, ), LLaVA (visualllama, ), and MiniGPT-4 (minigpt, ), show competitive performance but are typically limited to input resolutions of 2K-4K. To overcome this limitation, LLaVA-Next (llava-next, ) processes images in patches and connects them via global tokens, Monkey (monkey, ) and LLaVA-UHD (llava-uhd, ) compress patches to reduce redundancy, Cambrian (cambrian, ) uses learnable queries for multi-scale interaction, and SliME (slime, ) applies dual compression to preserve global context. In the RS domain, several MLLMs have been developed. For example, GeoChat (geochat, ) enables multi-task RS dialogue based on the LLaVA-1.5 framework (llava15, ). EarthGPT (earthgpt, ) unifies multisensor interpretation tasks by converting RS annotations into question-answer pairs. Meanwhile, LHRS-Bot (lhrsbot, ) improves vision-language understanding through multi-level alignment and curriculum learning. Nevertheless, both general-domain and RS MLLMs struggle to effectively handle UHR RS imagery. Universal MLLMs lack domain adaptation and are unable to meet the higher resolution requirements of real-world RS applications. Although RS-specific MLLMs incorporate domain knowledge, they still operate at significantly lower resolutions, i.e., less than 1K, highlighting the urgent need for a UHR ($\times8K$) and domain-aligned MLLMs in the RS field.
@@ 3 Dataset
To address this, following XLRS-Bench (xlrs-bench, ), we first manually annotated 12K UHR samples. Notably, existing MLLMs such as GPT-4o (gpt4o, ) either encounter memory overflow or generate low-quality outputs when processing UHR RS imagery. Therefore, we adopt a fully manual annotation approach in this work. Nevertheless, due to the high time and labor costs, scaling manual annotation of UHR data is impractical. To address this, we develop a semi-automated annotation pipeline that leverages GPT-4o (gpt4o, ) alongside existing RS detection and segmentation datasets, generating 100K medium-to-high-resolution (MHR, $\times2K$) image-text pairs. To bridge the distribution gap between MHR and UHR data, we further apply an influence-based data selection method built on the LESS framework xia2024less . Next, we detailed the procedure of constructing datasets.
@@
LLaVA-Next LLaVA-Next-2K_SFT llava-next 738K $\times512$ 44.6
[…]
ChatEarthNet ChatEarthNet 6k $\times256$
[…]
FIT-RS skysensegpt 1200k $\times512$
MMRS-1M earthgpt 308K $9×\times4$91
VRSBench-train vrsbench 85k $\times512$
@@
   - Multi-image questions 150
   - Average question length 128
   - Average answer length 22
   - Object category 18 (Coarse)
   - Object size (pixel ratio) 0.14%
[…]
@@
   - Average question length 172
   - Average answer length 41
   - Object category 60
   - Object size (pixel ratio) 1.02%
[…]
@@ 4.1 Overwhelming Background Tokens Hinder MLLM Fine-Tuning on RS Data
Figure 4: Examples of tokens and their positions that the logitlogit lenslens yields in the late layers.
In order to learn about a holistic recognition for the background ratio in UHR RS scenes, we conduct preliminary experiments, where the LLaVA-1.5 llava15 is employed. Specifically, we randomly selected 100 images from the XLRS-Bench (xlrs-bench, ), covering natural backgrounds such as sea surfaces, forests, and fields, to provide a diverse set of test samples for the evaluation of background redundancy. Then, we extract $\times64$ small images through non-overlapped sliding windows, and sequentially input each into the model. We evaluated the background rate of the UHR RS scene using two methods, following the previous work (ablation, ): Generative Description: We prompt the model with “Describe the image” and analyze the generated description to extract key information. Binary Polling: We ask the model “Is this image mainly background?” where the background is defined as low-information natural areas (e.g., sea, deserts, vegetation), excluding artificial structures (e.g. buildings, roads, urban areas).
@@
To better understand the semantic information of visual tokens, we use the logitlogit lenslens technique (LogitLens, ). For each layer, we decode the activation at each token position via unembedding. Details are shown in the appendix. With the logitlogit lenslens technique, we found that in large background areas, only a small portion of the information in the visual tokens is effectively mapped to specific semantic scenes, while most of the background tokens do not align with clear semantic representations (Fig. 4(b)).
@@
Then, we feed the whole UHR RS imagery into MLLM, where the LLaVA-Next-2K longva , a variant of LLaVA-Next llava-next , is adopted to support higher resolution as far as possible. Initially, we directly perform the inference on XLRS-Bench xlrs-bench , whose images have an average resolution of $\times8K$. Naturally, the resulting long visual token sequences lead to out-of-memory (OOM) errors, aligning with the first challenge presented in Section 1.
@@
To address the OOM issue, we increase the pooling layer size from the default $\times2$ to $\times6$, allowing LLaVA-Next-2K to process $\times9K$ resolution, called LLaVA-Next-9K. To further reduce redundant background information, we use a background-aware token selection strategy. By calculating semantic similarity between visual tokens and typical background terms (e.g., ocean, desert), we score and rank tokens, then discard the top 50% most redundant ones. The process involves: (1) extracting embedding vectors for typical background terms; (2) calculating each token’s highest similarity score with these embeddings as its background score; and (3) ranking tokens by score and removing the top 50%. This approach reduce the background token count. The result is referred to as LLaVA-Next-9K-Pruning.
@@ 4.2 Scarce Object Tokens Drive MLLM Fine-Tuning on RS Data
We conduct ablation experiments to test whether the foreground information is concentrated in specific visual tokens. By removing selected tokens and observing the drop in recognition performance, we assess their importance. The images are sourced from XLRS-Bench, with two pre-processing steps applied to ensure reliability: High signal-to-noise sub-images and Hallucination control. Details are shown in the appendix. Following this procedure, the final dataset includes 1,189 VQA pairs. Fig. 6 provides an overview of our ablation. Details are shown in the appendix.
@@
Tab. 3 shows that ablating object tokens significantly hinders the model’s ability to recognize targets. Generative decrease and VQA decrease experiments are following the previous work ablation . A larger percentage drop in performance indicates a greater impact of the ablation, meaning the model is more likely to answer incorrectly, thus suggesting that the ablated tokens contain more localized object information. Notably, with a comparable number of ablated tokens, removing object tokens consistently results in a larger performance degradation than random ablation, highlighting the precise localization of object-specific information. logitlogit lenslens analysis LogitLens further reveals that visual tokens for scarce objects efficiently converge to accurate semantic representations (Fig. 4 (a)). Collectively, these findings underscore a key insight in RS: object tokens not only encode critical information but also align closely with actual visual features.
@@ 4.3 Background Token Pruning and Anchored Token Selection
where $n$ is the number of remaining image tokens, $d$ is the dimension of hidden states, and ${\bm{W}}_{Q},{\bm{W}}_{V}$ are the query and key projection matrices of this encoder layer. Note that we utilize the attention map at the second-to-last layer of the visual encoder ($i.e.$, the output layer of CLIP-ViT in LLaVA-Next (llava-next, )). Given a compression ratio $r$, we calculate the final number of tokens to retain as $R=n\times r$. Ultimately, we select the top $R$ tokens with the highest attention scores from the attention map ${\bm{a}}_{[\texttt{CLS}]}$, obtaining the final retained tokens $Z_{v}^{\prime}\in\mathbb{R}^{R\times d}$.
@@ 5 Experiments
5.1 Main Results
Experimental Setup. We use XLRS-Bench xlrs-bench for evaluation. The MLLMs evaluated on XLRS-Bench are grouped into three categories: (a) open-source MLLMs; (b) closed-source MLLMs and (c) the specialized RS model. For fair comparison, we used a zero-shot setting with uniform prompts for all MLLMs, including our work. The appendix details the architecture and parameter sizes of each open-source MLLMs, and includes additional results across various settings. Except for GeoChat which was evaluated using its native framework, all other models were evaluated using LMMs-Eval zhang2024lmmsevalrealitycheckevaluation ; lmms_eval2024 . Following XLRS-Bench xlrs-bench , we evaluated the accuracy and reported of L-1 dimension for the VQA task, with L-3 and L-4 results available in the appendix.
@@ 5.1 Main Results
Main Results. After fine-tuning on SuperRS-VQA and HighRS-VQA, our GeoLLaVA-8K delivers outstanding performance across various evaluation tasks. It not only outperforms domain-specific models but also surpasses all existing open- and closed-source models, including the latest Qwen2.5 and InternVL3. Remarkably, with just 7B parameters, GeoLLaVA-8K even outperforms Qwen2.5-VL-72B, the largest and best-performing open-source MLLMs. This impressive gain stems from the high-quality dataset and the targeted compression strategy designed for the low semantic density of RS imagery.
5.2 Further Analyses
Table 5: FLOPs and latency of GeoLLaVA-8K under different compression ratios.
Compression Tokens/Grid Visual Tokens TFLOPs Latency (s/img) Avg.
[…]
$\times 24$ 14.0k 198.06 2.17 51.5
[…]
Effect of High-Resolution Data vs. Token Optimization Strategies We conducted an ablation study to analyze the effects of high-resolution data and token optimization strategies. As shown in Table 6, using high-resolution datasets (SuperRS-VQA and HighRS-VQA) already improves performance over the baseline. When combined with Background Token Pruning (BTP) and Anchored Token Selection (ATS), the model achieves the best average accuracy of 51.5%. This shows that while high-resolution data enhances visual understanding, most performance gains come from token optimization, which focuses attention on semantically important regions.
Table 6: Ablation study on the effect of high-resolution datasets and token optimization strategies.
Subtasks OC RC OLUC RLUC OCC OCL OMS OSR AD ECR RP RCCD CCR Avg.
LLaVA-Next-2k (Baseline) 28.3 47.0 3.0 62.0 42.0 34.1 66.7 27.2 73.0 75.0 34.0 48.3 40.0 44.7
+ High-Resolution Dataset 25.9 41.0 50.0 68.0 40.4 28.5 64.9 33.4 58.0 77.0 57.0 45.0 50.0 49.1
+ Dataset + BTP(OOM) - - - - - - - - - - - - - OOM
+ Dataset + BTP + ATS (GeoLLaVA-8K) 26.7 38.0 49.0 69.0 41.6 31.6 65.0 35.0 67.0 78.0 66.0 50.0 52.0 51.5
[…]
Small Object Preservation and CLS Attention under Token Compression. We analyzed the impact of token compression and CLS-based attention on small-object scenes using a subset of XLRS-Bench defined by COCO’s area criteria. Small objects were defined as occupying less than 5% of the image area, and categorized into three levels: Extremely Small (<0.1%), Very Small (0.1%-2.0%), and Normal Small (2.0%-5.0%). A dedicated subset of 1,570 samples covering six tasks (classification, color, motion state, counting, spatial relations) was curated. As shown in Table 7, GeoLLaVA-8K outperformed LLaVA-Next across all tasks, with notable improvements in spatial reasoning, counting, and motion state. These results show that token optimization effectively preserves small-region semantics and reduces information loss. The CLS attention mechanism generally focuses on task-relevant small objects but occasionally favors a single dominant region, suggesting room for further improvement in multi-object focus and attention balance.
Table 7: Performance comparison on the small-object subset of XLRS-Bench.
Dataset Object Classification Object Color Object Motion State Regional Counting Overall Counting Object Spatial Relationship Overall
LLaVA-Next 38.6 29.4 62.0 35.0 22.5 20.0 34.1
GeoLLaVA-8K 45.3 34.4 70.0 40.0 32.5 35.0 40.3
Table 8: Generalization performance on LRS-VQA.
Method Accuracy (%) Evaluation Format
LLaVA-Next llava-next 55.07 MCQ
GeoLLaVA-8K (ours) 56.28 MCQ
Generalization on External Datasets To evaluate the generalization of GeoLLaVA-8K, we tested it on the new ultra-high-resolution benchmark LRS-VQA lrsvqa . Since LRS-VQA only provides open-ended QA annotations, we converted it to a multiple-choice format using an LLM gpt4o to generate three plausible distractors per question, followed by structural validation. GeoLLaVA-8K achieved 56.28% accuracy, outperforming LLaVA-Next llava-next (55.07%) by 1.21%, confirming strong generalization to new datasets and formats.